OpenStudy (anonymous):

Complex conjugate anyone? x-4j^2+3j=yj+2x(-j)

OpenStudy (anonymous):

\[\Large x-4(j)^2+3j=yj+2x(-j) \] Like that?

OpenStudy (anonymous):

I figured it out. It follows: \[x-4j^2+3j=yj+2xj^3\] \[x-4(-1)+3j=yj+2x(-j)\] \[(x+4)+3j=yj+2x(-j)\] \[(x+4)+3j=0+(y+2x)(-j)\] \[(x+4)+3j=0+(y-2x)j\] formula \[a+bj=c+dj\] X ends up being -4 so that it equals 0 and y end up being -5 so the quantitiy (y-2x) equals 3. You're welcome.